OpenStudy (anonymous):

Help please. What is 275,000=1500(1-(1+.04/12))^t / (.04/12)

OpenStudy (anonymous):

I know I need to do ln(183.333) / ln (1-(1+.04/12)) / (.04/12) but my calculator won't give me an answer

OpenStudy (aum):

Is it \[ 275,000 = 1500\left(1 - \left (1 + \frac{0.04}{12}\right )\right)^t \div \frac{0.04}{12} \]?

OpenStudy (anonymous):

It is (1-(1+.04/12))/(.04/12). I divided by 1500 on both sides

OpenStudy (aum):

What happened to exponent t? Use the Equation / Draw button so it is clear what is being raised and what is on top and bottom.

OpenStudy (anonymous):

The exponent t is just for the (1-(1+.04/12))/(.04/12) equation. This is the